OpenStudy (anonymous):

What's a rational and irrational number between 7.7 and 7.9

OpenStudy (anonymous):

\[7\frac{ 8 }{ 10 }\] or 7.8

OpenStudy (jane11509):

Yea

OpenStudy (anonymous):

7.797979797979 is irrational.

OpenStudy (jane11509):

Yup

OpenStudy (anonymous):

It can't be expressed as a fraction because it has repeating decimals. It would actually be 7.79 with a line on top of 79.

OpenStudy (anonymous):

no ,you are wrong. let x=7.797979.... 100 x=779.7979.... subtract 100 x- 1x=772 99 x=772 x=772/99 which is rational.

OpenStudy (anonymous):

oki doke

OpenStudy (anonymous):

\[(7.7)^2=59.29\] \[(7.9)^2=62.41\] \[\sqrt{60},\sqrt{61}\] actually there are infinite irrational numbers between 7.7 and 7.9